### v2

# TEASMA: A Practical Approach for Test Assessment of Deep Neural Networks using Mutation Analysis

Abstract
Successful deployment of Deep Neural Networks (DNNs), particularly in safety-critical systems, requires their validation with an adequate test set to ensure a sufficient degree of confidence in test outcomes. Mutation analysis, a well-established technique for measuring test adequacy in traditional software, has been adapted to DNNs in recent years. This technique is based on generating mutants that ideally aim to be representative of actual faults and thus can be used for test adequacy assessment. In this paper, we investigate for the first time whether and how mutation operators that directly modify the trained DNN model (i.e., post-training operators) can be used for reliably assessing the test inputs of DNNs. Our results show that these operators, though they do not aim to represent realistic faults, exhibit strong, non-linear relationships with faults. Inspired by this finding and considering the significant computational advantage of post-training operators compared to the operators that modify the training data or program (i.e., pre-training operators), we propose and evaluate TEASMA, an approach based on post-training mutation for assessing the adequacy of DNN’s test sets. In practice, TEASMA allows engineers to decide whether they will be able to trust test results and thus validate the DNN before its deployment. Based on a DNN model’s training set, TEASMA provides a methodology to build accurate DNN-specific prediction models of the Fault Detection Rate (FDR) of a test set from its mutation score, thus enabling its assessment. Our large empirical evaluation, across multiple DNN models, shows that predicted FDR values have a strong linear correlation ($R^{2}\geq 0.94$) with actual values. Consequently, empirical evidence suggests that TEASMA provides a reliable basis for confidently deciding whether to trust test results or improve the test set of a DNN model.
Index Terms:
Deep Neural Network, Test Assessment, Mutation Analysis.
I Introduction
Effective testing of Deep Neural Networks (DNN) is essential to ensure they can be trusted, particularly as they are increasingly deployed in safety-critical systems. However, to rely on testing outcomes, it is crucial to assess the adequacy of the test set with which the DNN is to be validated. In the context of DNN testing, several techniques have been proposed for test adequacy assessment and are mostly relying on coverage criteria [1, 2, 3, 4, 5], inspired by practice in testing traditional software. However, these coverage criteria have shown no significant correlation with the number of mispredicted inputs or detected faults in a test set [6, 7, 5], thus putting into question whether they are reliable criteria for the test assessment of DNN models.
Mutation analysis, one of the most used techniques for test assessment in traditional software [8, 9], has been adapted to DNN models in recent years. DNN-specific Mutation Operators (MO) have been proposed [10, 11, 12, 13] since traditional MOs are not directly applicable to DNNs [14]. Indeed, developing a DNN model entails applying a training process relying on both a training program and a training set. Therefore, researchers have proposed MOs that fall into two primary categories: post-training MOs [11, 10], which mutate the DNN model obtained after the training process, and pre-training MOs [11, 13], which mutate either the training program or the training set and use these mutated versions in the training process to generate mutated models.
Post-training MOs were introduced by mutation testing frameworks for DNN models: MuNN [10] and DeepMutation [11]. Although DeepMutation introduces both the pre- and post-training MOs, its authors later introduced DeepMutation++ [12], a mutation testing tool implementing only the post-training MOs. Post-training MOs offer a practical advantage by not requiring retraining for mutant generation, resulting in vastly reduced computational costs compared to pre-training MOs. However, they directly modify a DNN model and thus are criticized for not realistically representing real faults. Hembatova et al. [13] further expanded the concept of pre-training MOs by investigating real faults in DNN models [14] and proposing corresponding, more realistic pre-training MOs. They also introduced a statistical approach to determine whether a mutant generated by pre-training MOs is killed [14]. However, this approach requires generating many instances of the original DNN model and each mutated model by repeating the training process many times for each model. Such an expensive procedure makes the cost of mutation testing impractical in many situations.
The use of mutation analysis in traditional software is motivated by the strong association between mutants and faults reported in the literature [15, 16, 8]. Regarding DNNs, the ability of MOs to generate killable and non-trivial mutants has been thoroughly investigated [11, 13]. However, the association between the mutants and DNN faults, and thus the ability to assess a test set based on its Mutation Score (MS), has never been studied. In this paper, we address this gap by investigating such an association for mutants generated from post-training MOs, motivated by their much lower cost compared to pre-training MOs. For this purpose we investigate, across a number of models and input subsets from training sets, the relationship between MS and Fault Detection Rates (FDR) considering various MS definitions proposed in the literature [11, 12]. We perform our experiments using post-training MOs implemented by DeepMutation++, a State-of-The-Art (SoTA) mutation testing tool for Feed-Forward Neural Networks (FNNs). We show that there are strong, non-linear relationships between FDR and MS, suggesting that mutants generated by post-training MOs could be used as fault indicators.
We also show that, despite the strong relationships between FDR and MS, their shape is not only highly non-linear but widely varies across subjects. This observation indicates that, in the context of DNN models, MS calculated based on post-training MOs is not a surrogate measure for a test set’s fault detection capability and thus cannot be used directly for test assessment. These findings motivate us to propose a new methodology that leverages MS to build an accurate and specific FDR prediction model for each DNN. To this end, we propose TEASMA, a practical solution designed to assess the adequacy of test sets for DNN models before deployment through mutation analysis based on post-training MOs. Following TEASMA, engineers can develop a specific FDR prediction model, along with Prediction Intervals (PI), for each DNN model based on its training set. This prediction model can then be used by engineers to predict a test set’s FDR based on its MS and decide, considering the prediction’s PI, whether a test set is adequate for validation in their context. Developing TEASMA is further motivated by the costly and time-consuming process of preparing large test sets, whose labeling usually requires domain experts. Using TEASMA, a test set is assessed before labeling and execution, solely based on the training set used for training the DNN. Test inputs can then be labeled and executed once the test set is deemed adequate, thus preventing the waste of testing resources. Further, if sufficiently fast and when MS is shown to accurately predict FDR, TEASMA could be used to automatically guide test selection.
We evaluated TEASMA on five widely used DNN models and four image recognition input sets. We compared, for a large number of test subsets, their predicted FDR with the actual one and we could observe, in all cases, a very high linear correlation ($R^{2}\geq 0.94$ ) between the two, with a slope close to 1 for the regression line. We also show that our prediction models have narrow PIs, implying that the predicted FDR can confidently be used as a basis for deciding about the adequacy of a test set.
To further justify our choice of using post-training MOs in TEASMA, we replicated our experiments using pre-training MOs implemented by DeepCrime, considering only one of our smaller DNNs to maintain the computations at a reasonable level and obtain a conservative ballpark estimate of the cost of using pre-training MOs. Our results demonstrate that performing mutation analysis, using the same inputs based on pre-training MOs and the statistical MS calculation of DeepCrime, incurs a computational time approximately 153 times greater than that required for using DeepMutation++’s post-training MOs and MS calculation approach. To investigate whether such a substantially higher cost would be justified by a stronger association between mutants and faults, we compared the correlation between FDR and MS computed based on these two tools. The findings show that using MS based on pre-training MOs does not yield any advantage over post-training MOs. These results show that conducting mutation analysis using post-training MOs not only requires significantly less computational resources compared to pre-training MOs but also leads to a MS that can serve as a reliable basis for predicting FDR. This further confirms our choice of relying on post-training MOs in TEASMA.
To summarize, the key contributions of this paper are as follows:
  • An empirical analysis of SoTA post-training MOs showing that these mutants have a strong relationship with faults in DNNs, though the shape of the relationship tends to be highly non-linear and varies widely across DNNs. Results thus confirm early suspicions that such MOs cannot directly be used as surrogates for faults and as such for test assessments.
  • TEASMA, a practical approach utilizing post-training MOs for assessing a test set before validation and deployment of a DNN. It is based on regression analysis using its training set, with the objective to build a predictor model for FDR, which can then be used as a basis for assessing the adequacy of test sets.
  • An empirical evaluation of TEASMA on five widely used DNN models and four image input sets, demonstrating its ability to accurately predict a test set’s FDR based on its MS, relying on a DNN-specific prediction model built during training.
  • An empirical analysis of the cost of applying pre-training MOs in comparison to post-training ones, to further justify our choice to rely on the latter.
The remainder of this paper is structured as follows: Section II provides background and discussions of the related work directly relevant to our research questions. Section III presents our proposed approach for test adequacy assessment. Section IV describes the experiments we performed to assess our approach. Section V presents and discusses the results for each research question. Section VI discusses the cost of pre-training MOs, and Section VII concludes the paper.
II Background and Related Work
This section is divided into four parts. In the first part, we briefly describe the proposed mutant generation processes for DNNs. The second part summarizes different ways of identifying killed mutants and calculating MS. In the third part, we discuss the association between mutants and faults and point out the challenge of investigating such association in DNNs. The last part describes a SoTA approach that can be leveraged to overcome this challenge and perform our experiments.
II-A Mutant Generation
Applying mutation testing to DNN-based systems requires defining specific MOs for DNN models [14]. The proposed MOs can be broadly classified into two primary categories, namely pre-training [11, 13] and post-training [11, 10, 12] operators, as illustrated in Figure 1. Pre-training MOs, also called source-level MOs, slightly modify the original version of either the program or the input set that has been used to train the original DNN model under test. These modified versions are then used by the training process to generate mutated models (i.e., mutants). Mutation testing using pre-training MOs is computationally expensive since it entails repeating the training process for generating each mutant. On the other hand, post-training MOs, also referred to as model-level MOs, modify the already trained original model, eliminating the need for additional training to generate mutants.
Figure 1: Pre-training and post-training mutation operators
Ma et al. [11] proposed DeepMutation, the first mutation testing framework for DNNs that includes both pre- and post-training MOs. Pre-training operators are classified into data-level and program-level operators whereas post-training operators are classified into weight-level, neuron-level, and layer-level operators, depending on the specific component of the DNN model that they modify. The DeepMutation++ tool [12] was later introduced by the same authors implementing only the post-training operators. Shen et al. [10] proposed the MuNN framework, including five post-training operators that modify neurons, activation function, bias values, and weight values of a DNN model. Since post-training MOs directly modify the DNN model, not all of them are applicable to any DNN model. For example, applying the Layer Deactivation ($LD$) operator of DeepMutation [11] on some layers may break the model’s structure. Therefore, DeepMutation restricts the application of this operator to only those layers whose input and output shapes are consistent, if the model has such layers.
Humbatova et al. [13] proposed DeepCrime, a mutation testing tool that includes 24 pre-training mutation operators inspired by real faults in deep learning systems [17, 18, 19]. The main goal was to obtain more realistic mutation operators. They also proposed a statistical approach [14] for identifying killed mutants that requires multiple instances of each mutant. Therefore, DeepCrime repeats the training process on the original and modified versions of the training set and program many times, resulting in many instances of the original model and each mutant. This further increases the cost of mutation testing, requiring not only to repeat the training process and generating more mutants but also executing each test set on all mutant instances.
Most of the MOs proposed for DNNs [11, 13, 10] have parameters that define the scale of modifications by that MO and need to be configured before their application. For example, operators that modify training sets could be configured by defining the percentage of inputs to be modified. Therefore, each MO can be applied to the DNN with multiple configurations, some of them resulting in generating mutants with undesirable characteristics: equivalent, trivial, and redundant. It is essential to identify and filter such mutants since they not only increase the overall execution time of mutation analysis but may also affect the MS. Different approaches have been proposed to address this issue [11, 12, 13]. In the DeepMutation++ tool, mutants with prediction accuracy lower than a specified threshold are eliminated. Then, for each input set, mutants with high error rates (i.e., killed by most of the inputs), are filtered out to prevent inflating the MS with easy-to-kill mutants. The DeepMutation++ sets the accuracy threshold to be 90% of the original model’s accuracy, and the error rate threshold to be 20% of the input set size. Humbatova et al. [13] defined killability and triviality scores for MOs in DeepCrime based on the training set. Redundant mutants are also identified and filtered out based on the subsumption relation between them [13]. Subsequently, to identify and exclude equivalent (non-killable), trivial, and redundant mutants, the entire training set is executed against all mutants generated by various configurations of each MO.
II-B Mutation Score Calculation
Similar to traditional software, killed DNN mutants can be identified by comparing the output of a mutant $M_{i}$ with the original model $N$ and the corresponding MS can be calculated as:
$StandardMS(T,M)=\frac{\sum_{M_{i}\in M}killed(T,M_{i})}{\left|M\right|}$ (1)
where $M$ is the set of mutants, $M_{i}\in M$ is $killed$ by test set $T$ ($killed(T,M_{i})=1$) if there is at least one input $t\in T$ that is correctly classified by $N$ and is not correctly classified by mutant $M_{i}$. DeepMutation [11], however, calculates MS in a different way. For a classification problem with $k$ classes $C=\{C_{1},...,C_{k}\}$, a test input $t\in T$ kills class $C_{j}\in C$ of mutant $M_{i}\in M$ if $t$ is correctly classified as $C_{j}$ by the original model $N$ and is not classified as $C_{j}$ by mutant $M_{i}$. Accordingly, MS is calculated for a test set $T$ as:
$DeepMutationMS(T,M)=\frac{\sum_{M_{i}\in M}killedClasses(T,M_{i})}{\left|M\right|\times\left|C\right|}$ (2)
where $killedClasses(T,M_{i})$ is the number of classes of mutant $M_{i}$ killed by inputs in $T$ [11]. The difference when calculating MS for the same test set using $StandardMS$ based on Equation (1) and $DeepMutationMS$ based on Equation (2) can be very large. For instance, consider a test set $TS$ from the input set of handwritten digits (MNIST). Each input $t\in TS$ is an image belonging to a class $C_{j}$. Let the test set $TS$ be represented by its correct input labels $label(TS)=\{3,5,3,7\}$, including two different images from the same class representing digit 3 but also two images from classes representing digits 5 and 7. Further, consider that all images are correctly predicted by $N$. If we have a set of mutants $M=\{M_{1},M_{2}\}$, where the first two test inputs of $TS$ are mispredicted by $M_{1}$ and all test inputs are mispredicted by $M_{2}$, then $DeepMutationMS(TS,M)=\frac{2+3}{2\times 10}=0.25$ while $StandardMS(TS,M)=1$. The score calculated by $StandardMS$ for a test set $T$ is higher since killing a mutant $M_{i}$ with $T$ requires only one test input $t\in T$ to be mispredicted by $M_{i}$.
Hu et al. [12] defined a killing score metric for each test input $t$ as the proportion of mutants in $M$ whose output is different from $N$.
$KillingScore(t,M,N)=\frac{\left|M_{i}|M_{i}\in M\wedge N(t)\neq M_{i}(t)\right|}{\left|M\right|}$ (3)
Humbatova et al. [13] used this metric to calculate the MS of a test set $T$ on mutants generated by DeepMutation++ [12]:
$KSBasedMS(T,M)=\frac{\sum_{t\in T}KillingScore(t,M,N)}{\left|T\right|}$ (4)
However, calculating MS as the average $KillingScore$ of test inputs in a test set may not indicate the ability of the entire test set to kill mutants. A test set cannot achieve such a high MS unless many of its test inputs are able to kill most of the mutants. This is visible in results reported by Humbatova et al. [13] where the MS calculated by $KSBasedMS$ for entire test sets ranges from 0.059 to 0.33. Humbatova et al. [13] further proposed a statistical approach for identifying killed mutants when applying pre-training MOs in DeepCrime where the training process is repeated $n$ times, leading to $n$ instances of the original model and mutants being generated (Section II-A). This further increases the cost of mutant generation using pre-training MOs and makes them computationally challenging for practical use in our context.
Humbatova et al. [13] proposed a statistical approach for identifying killed mutants when applying pre-training MOs in DeepCrime where the training process is repeated $n$ times and $n$ instances of the original model $N=\{N_{1},...,N_{n}\}$ and each mutant $M_{i}=\{M_{i1},...,M_{in}\}$ are generated (Section II-A). A mutant $M_{i}$ is considered killed by DeepCrime if, for a given test set $T$, the difference between the accuracy of $N(T)$ and $M_{i}(T)$ is statistically significant:
$isKilled(N,M_{i},T)=\begin{cases}1,&\text{if $p-value<\alpha\quad\textrm{and}\quad effectSize\geq\beta$}\\ 0,&\text{otherwise}\end{cases}$
The $p-value$ is obtained by using the Generalized Linear Model (GLM) [20] with $\alpha=0.05$ and $effectSize$ is calculated using Cohen’s d [21] with $\beta=0.5$. The MS of a test set can be calculated directly using $isKilled$ as follows:
$isKilledBasedMS(T,M)=\frac{\sum_{M_{i}\in M}isKilled(N,M_{i},T)}{\left|M\right|}$ (6)
DeepCrime [13], however, defined MS differently for a test set $T$ on a set of mutants generated by a given mutation operator $MO_{i}$:
$OperatorMS(MO_{i},T)=\frac{\left|K(MO_{i},T)\cap K(MO_{i},TrSet)\right|}{\left|K(MO_{i},TrSet)\right|}$
where $K(MO_{i},T)$ denotes those mutants generated by $MO_{i}$ that are killed by $T$, and $TrSet$, is the training input set used to generate the original model $N$. Only the intersection of mutants that are killed by both the test set and the training input set is considered in order to exclude equivalent mutants. Finally, the overall mutation score of $T$ on all mutants is determined as the average of mutation scores calculated for each individual operator:
$DeepCrimeMS(T,M)=\frac{\sum_{MO_{i}\in MO}OperatorMS(MO_{i},T)}{\left|MO\right|}$
This MS calculation captures the ability of a test set to kill mutants generated by all MOs and a test input subset can only achieve a high MS when it is able to kill a large proportion of mutants generated by each MO. It is also important to note that since the killed mutants are determined by comparing the accuracy of the original model and mutants based on Equation(5), the correct label of each test input must be known. Therefore, we cannot calculate the MS of an unlabeled test set before deployment with this approach.
To be able to use the pre-training MOs of DeepCrime and still calculate the MS of a test set before deployment, we can rely on comparing the output of a mutant with the original model and calculate MS in the same way as DeepMutation ($DeepMutationMS$ in Equation (2). This way we can also fairly compare the MS of a given test set on mutants generated by DeepCrime pre-training operators with DeepMutation post-training operators. For this purpose, we need to determine the output of each mutant and the original model. Since DeepCrime generates $n$ instances of the original model and each mutant, one way to determine the output is to rely on majority voting by considering the most voted output by all $n$ instances. Consequently, the output of a mutant $M_{i}$ (or the original model $N$) for a given test input $t$ can be identified as:
$M_{i}(t)=mode\{M_{i1}(t),M_{i2}(t),...,M_{in}(t)\}$ (9)
Mutation analysis involves identifying test sets with high MS based on the key assumption that MS captures a test set’s capability to detect faults. In DNNs, however, various MS formulas lead to different results and which test sets lead to high MS varies according to which formula is used. It is evident that more test sets can achieve a high MS using $StandardMS$ (Equation (1)) than when using $DeepMutationMS$ (Equation (2)). We also expect that not many test sets can achieve a high MS using $KSBasedMS$ (Equation (4)). Therefore, because we cannot a priori determine which MS formula satisfies our assumption to the best extent, it is essential to consider all of them when experimenting with mutation analysis for DNNs.
II-C Mutation Analysis Cost
The high computational cost of mutation analysis is its main drawback in practice. This cost is mainly due to performing two processes. Initially, one must utilize MOs to generate mutants. The second process involves executing a large number of test inputs against mutants, followed by the calculation of their MS. A precise calculation of MS usually necessitates filtering out the equivalent, trivial, or redundant mutants.
The cost issue is even more critical in the case of DNNs considering large input sets, which is usual in practice. In this context, the substantial computational overhead of mutant generation using pre-training MOs, due to the need to perform training on each mutant, is a significant drawback when compared with post-training MOs. This is particularly true for DeepCrime which involves generating many instances of each mutant, even if only the most effective configurations [14] are considered for each MO to create non-equivalent and non-trivial mutants. Subsequently, computing MS is even more time-consuming as each input is run against all instances of each mutant. In order to characterize the differences in computational cost between mutation analysis employing pre- and post-training MOs, we report in Table I a formula capturing the average cost of each process as a function of parameters explained below. This is based on the assumption we use two SoTA mutation testing tools for DNNs, namely DeepMutation++ and DeepCrime.
TABLE I: Average cost function of performing mutation analysis using pre- or post-training MOs
Tool Mutant generation MS calculation
DeepMutation++ $mutateTime(N)\times\left|M_{G}\right|$ $execTime(T)\times(\left|M_{F}\right|+1)$
DeepCrime $mutateTime(TrSet,TrP)\times\left|M_{G}\right|$ $execTime(T)\times(\left|M_{F}\right|+1)\times n$
$+trainTime(M)\times\left|M_{G}\right|\times n$
First, $mutateTime(N)$ represents the average time to create a mutated version of the original model $N$. Similarly, $mutateTime(TrSet,TrP)$ denotes the average time to create a mutated version of either the original training set $TrSet$ or the training program $TrP$, the $trainTime(M)$ denotes the average time to train a DNN model, and the $execTime(T)$ represents the average time to execute a test set $T$ on either the original or a mutated DNN model. Additionally, $\left|M_{G}\right|$ is the number of generated mutants, and $\left|M_{F}\right|$ is the number of mutants after filtering out those considered equivalent, trivial, or redundant. The primary difference in mutant generation cost between any pre-training and post-training mutation testing tool stems from the retraining involved for each pre-training mutant. Given that DeepCrime generates $n$ instances for each mutant, the retraining process is performed $n$ times for every mutant. Consequently, the computational time required for calculating MS based on mutants generated by DeepCrime is $n$ times greater than that of DeepMutation++. Given the necessity of computing MS for numerous test sets during mutation analysis, the execution of these test sets results in a significant difference in computational cost between the two tools.
Considering such substantial differences, it is essential to investigate if the more practical post-training MOs are also able to generate mutants associated with faults that can support our objective, that is assessing the adequacy of test sets. In this paper, we perform such an investigation and propose TEASMA, a practical approach for test set adequacy assessment in DNNs based on mutation analysis using post-training MOs. Further, to support our analytical arguments above, we report on the cost of mutation analysis deploying pre-training MOs for one of our simplest DNNs and input sets in SectionVII. Results show that relying on DeepCrime’s MOs and MS calculation formula (Equation (8)) took about 153 times the amount of computation time of DeepMutation++. This ratio should vary across subjects and is logically expected to be higher for larger DNNs and input sets. Nevertheless, our observations confirm the expected high cost of using pre-training mutants.
Figure 2: The process of test adequacy assessment with TEASMA
II-D Mutants and Faults
In traditional software, MOs are defined to introduce simple syntactical changes representing real mistakes often made by programmers into the source code of a program. Similarly, DeepCrime defined a set of pre-training MOs based on an analysis of real faults [17, 18, 19] related to the training set or program used to train a DNN model. However, the high cost of generating mutants using pre-training MOs hinders their application in practice. Post-training MOs, in contrast, apply modifications directly to the original model but have been criticized for performing modifications that are not realistic [22]. Nevertheless, for mutation analysis to be effective in assessing the adequacy of test sets, regardless of how mutation operators are defined, it is important to generate mutants whose corresponding MS is, for a test set, a good predictor of fault detection. Therefore, considering the strong practical advantages of post-training MOs, despite their presumed lack of realism, they should be carefully investigated to build predictors of fault detection before being dismissed as inadequate.
The use of mutation analysis for assessing the quality of test sets in traditional software is supported by the observed association between mutants and real faults, as reported in the literature [15, 16, 8]. This association has been investigated by performing correlation analysisbetween the MS of a test set and its capability to detect faults. Regression analysis has also been applied for modeling this relationship [15, 8]. In DNNs, however, the association between mutants and faults has never been investigated. The main challenge in such investigation is that, in contrast to traditional software where it is possible to identify real faults by isolating the faulty statement that caused a failure, identifying faults in DNNs is not a straightforward task due to the complexity and non-linear nature of the DNNs. However, a recent approach proposed by Aghababaeyan et al. [6] for estimating faults is a possible, practical, and automated solution. We describe next how we rely on this approach to investigate the relationship between mutants and faults, and how MS can help predict FDR.
II-E Fault Estimation
The number of mispredicted test inputs is a well-known metric in evaluating DNN testing approaches [1, 23, 24]. However, the number of mispredictions is not a good indicator of the number of faults and hence it is not an adequate metric to perform mutation analysis. Indeed, a single fault within a DNN model can make the model mispredict multiple test inputs. Subsequently, a test set that identifies a considerable number of mispredictions may actually detect only a few underlying faults, whereas another test set with the same number of mispredictions could detect a higher number of distinct faults. Therefore, assessing test sets based on mispredictions and thus investigating the relationship between the number of killed mutants and mispredictions can be misleading.
Consequently, we rely on a recent approach introduced by Aghababaeyan et al. [6] for estimating faults in a DNN model, that are defined as distinct root causes of mispredictions [25]. Aghababaeyan et al. [6] cluster mispredicted inputs that exhibit similar features in three key steps: feature extraction, dimensionality reduction, and density-based clustering. Initially, VGG-16 [26] is employed to extract the feature matrix from the mispredicted inputs, serving as the foundation for clustering. Subsequently, dimensionality reduction is applied to enhance clustering performance within the inherently high dimensional feature space. The HDBSCAN algorithm [27] is then utilized to cluster the mispredicted inputs based on the extracted features. Aghababaeyan et al. [6] empirically analyzed the resulting clusters and showed that inputs in the same cluster are mispredicted due to the same fault in the model and, conversely, inputs in different clusters are mispredicted due to distinct faults. Therefore, if a test set contains only one of the test inputs in a cluster, then it is able to detect the underlying fault of that cluster.
Suppose we identify a set of $k$ Clusters of Mispredicted Inputs $CMI=\{CMI_{1},...,CMI_{k}\}$, we can then assume that each cluster $CMI_{i}$ contains inputs that are mispredicted due to the same root cause (i.e., the underlying fault $F_{i}$). Therefore, we say that test set $T$ $detects$ a fault $F_{i}\in F$ if there is at least one test input $t\in T$ such that $t\in CMI_{i}$. Finally, we calculate the FDR for a test set $T$ as:
$FDR(T,F)=\frac{\sum_{F_{i}\in F}detect(T,F_{i})}{\left|F\right|}$ (10)
In the next section, we propose TEASMA, a practical solution for test set adequacy assessment using mutation analysis. The objective is to predict the expected FDR for a given test set and DNN, based on mutation and regression analysis performed during training. We should note that while TEASMA currently utilizes the fault estimation approach described in this section, it is flexible and can incorporate new fault identification approaches to improve its performance.
III The TEASMA Approach
In this section, we present TEASMA, a practical solution for assessing test sets for DNNs using mutation analysis. The basic idea is to rely on mutation and regression analysis based on the training set to accurately predict the FDR of test sets, which can then be used for evaluating their adequacy before proceeding with labeling and execution. We provide an overview of the main steps to apply TEASMA in Figure 2 and a detailed description of these steps in Algorithms1 and 2. The process to be followed by engineers in practice is illustrated in Figure 2 and includes three high-level steps:
1) Calculating MS and FDR for subsets of the training set.
In this step, only post-training MOs are used to generate mutants as described in Section II-A, since in our application context pre-training MOs are not a practical option. Faults of the DNN model are estimated using the misprediction clustering approach described in Section II-E based on the training input set. Then, a large number of subsets of varying sizes are sampled from the training set, and their MS and FDR are calculated to be used in the next step.
2) Performing regression analysis.
In this step, regression analysis between FDR and MS is performed to obtain the best regression model for predicting FDR based on MS. This prediction model is then used in the next step to assess the adequacy of test sets.
3) Assessing test set adequacy.
In this step, given a test set, its MS is calculated using the original DNN model and previously generated mutants. Then, the prediction model obtained from the previous step is used to predict the test set’s FDR ($\widehat{FDR}$) along with Prediction Intervals (PI).
The Algorithm 1 describes the first two steps implemented in TEASMA to create an accurate regression model for predicting FDR based on MS. As illustrated in Figure 2, the first two steps of TEASMA are applied entirely on the training set of the DNN. The algorithm starts by initializing an empty archive of subset samples $A$ (line 1). Then, to be able to perform mutation analysis, TEASMA is required to generate mutants (line 2) and identify faults in the DNN (line 3). TEASMA leverages DeepMutation++ to generate and select the set of mutants $M$ (Section III-A). It also relies on a SoTA approach to identify the set of distinct faults $F$ in the DNN model based on the model’s mispredicted inputs in the training set (Section III-B).
Input: $N$: the original DNN model under test $TrSet$: the labeled training input set of $N$ $SN$: number of samples per each sampling size
Output: $PM$: prediction model built for N
$sample\ archive\ A\xleftarrow{}\emptyset$   Generate post-training mutants $mutants\ M\xleftarrow{}GenerateMutants(N);$   Estimate faults based on the training set $faults\ F\xleftarrow{}EstimateFautls(N,TrSet);$   1) Calculate MS and FDR for training input subsets do
1       $size\xleftarrow{}UpdateSamplingSize(A);$ $samples\xleftarrow{}Sampling(TrSet,size,SN,type);$ foreach subset $sub_{i}$ in $samples$ do
2             $MS_{i}\xleftarrow{}MSCalc(sub_{i},M,N,MSFormula);$ $FDR_{i}\xleftarrow{}FDRCalc(sub_{i},F);$
3      $A\xleftarrow{}AddtoArchive(samples,MS,FDR);$
while $(0\leq FDR\leq 1\ is\ not\ covered)$  2) Perform regression analysis and select best model $regression\ models\ RM\xleftarrow{}\emptyset$ $RM\xleftarrow{}RegressionAnalysis(A);$ $PM\xleftarrow{}argmax(RM,R^{2});$ return $PM$
Algorithm 1 Building a specific prediction model for a DNN using its training set
The algorithm includes a main input sampling loop for selecting a large number of input subsets with various sizes from the training set of the original DNN model $TrSet$ and for calculating their MS and FDR (lines 4-11). The sampling is performed until the entire FDR range ($0\leq FDR\leq 1$) is covered. In each iteration, TEASMA updates the sampling size (line 5) with a larger or smaller size according to the minimum and maximum FDR achieved by the samples currently present in the archive $A$. At the first iteration, when the archive is empty, the $UpdateSamplingSize$ function starts with a small $size$. To be able to build a regression model, TEASMA requires multiple input subsets of the same size. Therefore, for each subset $size$, TEASMA samples $SN$ subsets from the training set (line 6). $SN$ is one of the input parameters of the algorithm and can be defined based on the size of the training set. For the reasons we describe later in Section III-C, TEASMA can sample input subsets either randomly or uniformly. After sampling input subsets, TEASMA calculates the MS and FDR of all subsets (lines 7-9). The MS of each subset can be computed based on one of the MS calculation approaches ($MSFormula$) described in Section II-B. In the following sections, we investigate the relationship between FDR and MS, as calculated using each MS calculation approach, to determine the best approach for building an accurate FDR prediction model. FDR is then computed based on the formula specified in Section II-E. Finally, the MS and FDR corresponding to all sampled subsets are added to the archive $A$ (line 10).
With a large number of input subsets of diverse sizes selected from the training set and their corresponding MS and FDR, TEASMA builds a set of regression models (lines 12 and 13). For reasons we will explain in Section III-D, TEASMA considers non-linear regression models including quadratic, exponential, and regression trees in addition to linear regression. To enable the engineer to identify the best regression model, TEASMA relies on common metrics for evaluating such models (Section III-D) including the coefficient of determination ($R^{2}$) to measure the goodness of fit. Using these metrics, engineers can determine for each DNN and training input set, the best regression model for predicting FDR based on MS. As a default, TEASMA selects the regression model with the highest goodness of fit ($R^{2}$) (line 14).
Input: $T$: an unlabeled test set to be assessed $N$: the original DNN model under test
Output: $\widehat{FDR}(T)$: predicted FDR of $T$ PI: prediction interval of $\widehat{FDR}(T)$
  Load mutants generated for N $M\xleftarrow{}LoadMutants(N)$   Load the prediction model built for N $PM\xleftarrow{}LoadPredictionModel(N)$   3) Assess test set T based on its MS $MS_{T}\xleftarrow{}MSCalculation(T,M,N,MSFormula);$ $[\widehat{FDR},PI]\xleftarrow{}Predict(PM,MS_{T})$ return [$\widehat{FDR}\ ,\ PI$]
Algorithm 2 Assessing test sets
Algorithm 2 describes the third step of TEASMA. As illustrated in Figure 2, this step is performed on a set of unlabeled test inputs before deployment. Given a test set $T$ for the DNN $N$ to assess, TEASMA starts by loading the same set of mutants $M$ generated in the first step (line 1) and the prediction model $PM$ built in the second step (line 2). To predict the FDR of a test set $T$, TEASMA first calculates MS (line 3) using the same MS calculation approach $MSFormula$ that has been used in the first step. Subsequently, the MS of the test set $T$ is used to predict its FDR based on the loaded prediction model $PM$ (line 4). TEASMA provides engineers not only with the predicted FDR of $T$ but also with its prediction interval $PI$ to enable them to decide on the adequacy of a test set considering the lower or higher bounds of the PI (Section III-E).
We argue that our proposed solution—assuming an accurate FDR prediction model can be built specifically for each DNN using its training set—is practical. Indeed, TEASMA relies on post-training MOs which are directly applied to the DNN model and do not entail any re-training, thus avoiding the main challenge with pre-training MOs. Algorithm 1 is executed once for a specific DNN to generate mutants and create the prediction model. Subsequently, Algorithm 2 can be employed multiple times to assess many test sets without incurring additional costs. In practice, such assessment helps guide the selection of test inputs and significantly reduces test costs by only labeling a test set once it is considered adequate. However, since we build the FDR prediction model using the training set and use it to determine the adequacy of a test set, we implicitly assume that the two sets have similar distributions. This is, however, a common assumption for the training of DNNs to be considered adequate.
Throughout the remainder of this section, we provide further details on key elements of TEASMA and its application to image classification input sets.
III-A Calculating Mutation Scores
To calculate MS precisely, in addition to configuring MOs correctly, we need to complement the mutant generation process by identifying and excluding mutants that are trivial or equivalent. TEASMA uses DeepMutation++ [12], the most recent SoTA mutation testing tool for DNNs that implements eight post-training MOs suitable for FNNs. Some of the implemented MOs need to be configured by setting values for their parameters. For example, the Neuron Effect Blocking (NEB) operator is configured by determining the number of neurons it affects. By default, the DeepMutation++ tool configures this operator to sample a predefined ratio of 1% of the total neurons in a DNN model and limits the total number of generated mutants to 50. TEASMA uses the default configuration for generating mutants. Furthermore, it performs the same procedure proposed by the DeepMutation++ tool and uses its default thresholds to filter out some of the generated mutants. Specifically, it excludes mutants when they yield less than 90% of the original model’s accuracy or mispredict more than 20% of the inputs correctly predicted by the original model, as reported in the original papers [11, 12]. Additionally, TEASMA performs an extra step to filter out equivalent mutants. The identification of equivalent mutants is achieved through a procedure similar to that of DeepCrime [13], involving the execution of the entire training set against the mutants. Mutants that cannot be killed by any training input, are considered equivalent and hence excluded. TEASMA can employ various approaches for MS calculation as explained in Section II-B. Consequently, in our experiment, we examine the correlation between FDR and MS calculated using each approach and empirically determine which approach results in the most accurate prediction model.
III-B Estimating Faults
TEASMA estimates faults in the DNN model using the approach described in Section II-E based on the mispredicted inputs of the training set. As we mentioned earlier, we implicitly assume that both the training set and the test set have a similar distribution. Except for transfer learning [28] where there is an expected distribution shift between the training set and the test set, this default assumption is correct and necessary for the training process to be considered adequate. Therefore, we can use the training set of the DNN model to identify faults of the DNN model and create its specific FDR prediction model. Note that, unlike test inputs, training inputs are labeled, and mispredicted inputs can be identified. To accomplish this, the entire training set is executed against the original DNN model, the mispredicted inputs are identified, and subsequently, the fault identification approach [6] is applied to these mispredicted inputs. Note that TEASMA builds a specific fault prediction model for each individual DNN based on its training set. Consequently, the performance of TEASMA can potentially be enhanced by integrating new fault identification approaches and creating more precise prediction models.
III-C Sampling Input Subsets
TEASMA, selects input subsets from the training set, calculates the MS and FDR of each subset, and builds a regression model. For this purpose, it requires a large number of input subsets that possess diverse MS and FDR values. Therefore, starting with sampling input subsets with a small size, TEASMA updates the sampling size and continues with larger or smaller sizes where needed, until the entire spectrum of FDR with minimum and maximum values ($0\leq FDR\leq 1$) is covered. To build an accurate prediction model, TEASMA requires the sampling of multiple subsets along the size range. The number of subsets sampled for each size is determined through an input parameter ($SN$) provided to TEASMA, which we set to 300 in our experiments. Sampling is performed with replacement, allowing for the potential inclusion of the same test input in multiple subset samples. This approach serves the purpose of including diverse testing inputs and variations of MS and FDR values across various samples. Since one of the MS formulas we use (Equation (2)) considers the number of images from each class $C_{i}$ in an input subset, TEASMA complements the conventional approach of creating randomly sampled subsets by creating a set of uniformly sampled subsets. Random subsets may contain any number of input images from each class $C_{i}$ but uniform subsets contain an equal number for each class $C_{i}$. Uniform subsets may achieve higher MS values than random subsets based on Equation (2). Therefore, we perform our experiments with both types of input subsets and empirically determine which one serves our purpose better.
III-D Regression Analysis
In this step, TEASMA uses all the sampled input subsets from the training set to perform regression analysis between MS and FDR. Given that post-training MOs may not be representative of real faults, we can expect non-linear relationships. In fact, as our subsequent results demonstrate, the shape of the relationship between MS and FDR is often highly non-linear and significantly varies across DNNs. Therefore, TEASMA considers quadratic and exponential regression models, as well as regression trees when no regression function fits well. Regression trees enable the optimal partition of the MS range into optimal sub-ranges with similar FDR values while avoiding overfitting. Note that the best model shapes may differ across DNNs because of differences in their architecture and training sets. Consequently, for each DNN model, the best shape for the regression model is determined empirically, prioritizing simpler linear models when they fare well. Regardless of its shape, the regression model must be sufficiently accurate to enable practically useful FDR predictions. To measure the goodness of fit for each regression model, TEASMA calculates the resulting coefficient of determination ($R^{2}$). TEASMA also computes two common measures for evaluating the accuracy of predictions, namely the Mean Magnitude of Relative Error (MMRE) and the Root Mean Square Error (RMSE) to assess how accurately FDR can be predicted. To obtain more realistic results, TEASMA performs a $K$-fold cross-validation procedure with $K=5$ and reports the average $R^{2}$, MMRE, and RMSE across all folds. Based on these metrics, engineers can determine for each DNN, the best regression model for predicting FDR based on MS.
III-E Assessing Test Sets
In the last step, when provided with a test set, TEASMA computes its MS using the original DNN and the previously generated mutants. Subsequently, it predicts the test set’s FDR ($\widehat{FDR}$), including PIs based on the prediction model obtained from the previous steps. Given a certain confidence level (e.g., 95% in our experiments), such intervals specify the range of potential actual FDR values for test sets achieving a specific MS value and thus provide engineers with valuable insights into the uncertainty of FDR predictions. Depending on the context, one can then decide how conservative should decision-making be by considering the lower or higher bounds of the PI. The $\widehat{FDR}$ of a test set, if deemed sufficiently accurate based on its PI, can then be used to assess its capability to detect faults and therefore whether, in a given context, it is adequate to validate the DNN.
IV Experimental Procedure
In this section, we describe the empirical evaluation of TEASMA, including the research questions we address, the subjects on which we conducted our evaluation and a comprehensive description of our experiments.
IV-A Research Questions
Our experiments are designed to answer the following research questions:
RQ1: Can an accurate regression model be built to explain FDR as a function of MS?
In this question, we investigate the output of the second step of TEASMA in Section III. We select a large number of input subsets from the training set, calculate their MS and FDR using the faults estimated in the original model as described in Section III-B, and investigate, for mutants derived from post-training MOs, if the MS of an input subset is significantly correlated with its FDR. In that case, we resort to regression analysis to model the relationship between MS and FDR and investigate if an accurate regression model can be built to predict a subset’s FDR from its MS. If that is the case, the predicted FDR can serve as a good basis on which to decide whether a test set provides sufficient confidence about the validation of the DNN before its deployment. Though such questions have been investigated for traditional software [15, 8], that has not been the case for DNNs.
RQ2: Can a test set’s FDR be accurately predicted from its MS using the regression model built on the training set?
In this question, we evaluate the last step of TEASMA in Section III and investigate if we can accurately predict a test set’s FDR using the regression model built in the previous step on the training set. For this purpose, we select a large number of input subsets from the test set and calculate their MS and FDR using the faults identified in the original model. Then, we use the regression model to predict FDR for each subset and compare it with the actual one, to assess whether, as expected, they have a linear relationship with a high correlation and a slope close to 1.
IV-B Subjects
TABLE II: Information about experimental subjects
ID Input Training Test Model Epochs Train Test Number of mispredictions Number of faults
set set set accuracy accuracy (training) (training)
S1 MNIST 60,000 10,000 LeNet5 12 96% 96% 2218 307
S2 Cifar-10 50,000 10,000 Conv-8 20 94% 85% 2507 354
S3 ResNet20 100 97% 89% 1437 191
S4 Fashion-MNIST 60,000 10,000 LeNet4 20 94% 90% 3579 512
S5 SVHN 73257 26032 LeNet5 20 94% 87% 4257 607
S6 VGG16 10 97% 93% 5469 590
We performed our study on different combinations of input sets and models. Since we rely on a SoTA fault estimation approach for DNNs that is tailored for image inputs, we selected a set of widely-used publicly available image input sets including MNIST [29], Fashion-MNIST [29], Cifar-10 [30], and SVHN [31]. Some of these input sets have also been used for evaluating SoTA mutation testing tools and operators [11, 10, 12, 13]. MNIST is a set of images of handwritten digits and Fashion-MNIST contains images of clothes that are associated with fashion and clothing items. Each one of MNIST and Fashion-MNIST represents a 10-class classification problem and consists of 70,000 grayscale images of a standardized 28x28 size. Cifar-10 is another widely-used input set that includes a collection of images from 10 different classes (e.g., cats, dogs, airplanes, cars). SVHN [31], a set of real-world house numbers collected by Google Street View, can be considered similar to MNIST since it includes images representing a sequence of digits from 10 classes of digits. The Cifar-10 and SVHN input sets contain 32x32 cropped colored images.
In our experiments, we trained five SoTA DNN models using the above input sets: LeNet5, LeNet4, ResNet20, VGG16, and an 8-layer Convolutional Neural Network (Conv-8). Since not all of the post-training MOs are applicable to all DNN models, as described in Section II-A, we selected DNN models with different internal architectures. All the feasible combinations of DNN models and input sets we experiment with are referred to as subjects. Overall, we perform our experiment on six subjects including a wide range of diverse image inputs and DNN architectures. Table II lists the details of each subject including the size of test and training datasets, the number of epochs used to train the model, the accuracy of the DNN model on the training and test input sets, and the number of mispredicted inputs and distinct faults identified based on the training set. Note that we reused the training and test sets as defined in the original sources [29, 31, 30].
IV-C Experiments
This section explains how we conduct our experiments to evaluate TEASMA and answer the research questions described in Section IV-A. To answer RQ1, we perform the first step of the TEASMA approach (Section III) as detailed in Algorithm 1 using the training input set of our subjects. As a result, we calculate FDR and MS for a large number of diverse input subsets sampled from the training set. Subsequently, we measure the correlation between FDR and MS using the non-parametric Spearman coefficient based on all subsets, since we cannot assume linearity. We also measure the correlation for mutants generated by each MO separately and then attempt to find subsets of operators leading to higher correlation. Before proceeding to the next step, we compare the highest correlation achieved by subsets of MOs with that obtained with all MOs. If the correlation of the former is higher than the latter, then we only consider mutants generated by the optimal subset with the highest correlation. Next, using all the training subsets, we proceed to the second step of TEASMA, regression analysis, relying on both linear and non-linear models. TEASMA performs a $K$-fold cross-validation procedure with $K=5$ and reports the average of $R^{2}$, MMRE, and RMSE across all folds for each regression model separately. The best regression model can be selected based on these metrics.
To summarize, the experiment that we conducted using the training set to answer RQ1 included the following steps:
  1. 1.
    Perform the first step of TEASMA and calculate MS and FDR for each input subset
  2. 2.
    Measure the Spearman correlation between FDR and MS
  3. 3.
    Perform the second step of TEASMA and conduct regression analysis between FDR and MS using linear and non-linear models (quadratic, exponential, and regression trees)
  4. 4.
    Evaluate regression models to identify the best one based on their $R^{2}$, MMRE, and RMSE calculated based on $K$-fold cross-validation
To address RQ2, we sample a large number of input subsets from the test set, calculate their MS using previously generated mutants, and leverage the previously built and selected regression model to predict FDR ($\widehat{FDR}(T)$). We should note that we determine killed mutants by comparing their output with that of the original model and thus the actual labels of test inputs are not required. To estimate the actual FDR ($ActualFDR$) of each test subset, we use the same fault clusters identified based on the training set as described in Section III-B and we assign each mispredicted test input to one of the clusters. Since the fault estimation approach (Section II-E) employs the HDBSCAN algorithm [27], a density-based hierarchical clustering method where each cluster is characterized by a number of core points, we assign each mispredicted test input to the cluster featuring a core point that is closest to the input. To ensure an accurate calculation of $ActualFDR$ and allow the test set to reach the maximum FDR value of 1, we exclusively take into account fault clusters that can be detected by the test set. To achieve this, we limit our denominator in Equation (10) to the number of fault clusters with at least one mispredicted test input assigned to them. Consequently, we conduct our experiments based on the faults detectable by the test set, so that our results are not affected by the varying quality of the provided test sets across subjects.
Subsequently, we are able to compare for each test subset, the predicted FDR with the actual one and investigate if the former is closely aligned with the latter and can thus be trusted to make decisions about the test set. Therefore, to answer RQ2, we went through the following steps:
  1. 1.
    Randomly sample a large number of input subsets of diverse sizes from the test set
  2. 2.
    Calculate MS and FDR ($ActualFDR$) for each test subset
  3. 3.
    Use previously selected prediction model to compute $\widehat{FDR}$ based on MS for each test subset
  4. 4.
    Based on all test subsets, measure the linear correlation ($R^{2}$) between $\widehat{FDR}$ and $ActualFDR$ and check the slope of the regression line
As described in Section II-B, different methods have been proposed to determine whether a mutant is killed and therefore to calculate MS. To fully address our research questions, we ran a comprehensive analysis by considering all the MS calculation variants described in Section II-B. As a result, we ran three variations of our experiments that only differ in the way they calculate MS: (E1) $DeepMutationMS$, (E2) $StandardMS$, and (E3) $KSBasedMS$.
Experiments on all subjects took about two days to run using a cloud computing environment provided by the Digital Research Alliance of Canada [32], on the Cedar cluster with a total of 1352 NVIDIA P100 Pascal GPUs with 8 to 16 GB of memory.
V Data Availability
The replication package for our studies will be shared upon the paper’s acceptance.
TABLE III: Spearman correlation coefficients computed between MS and FDR for random input subsets of the training set for each subject
Experiment Subjects
Variants S1 S2 S3 S4 S5 S6
E1 0.99 0.99 0.99 0.99 0.99 0.98
E2 0.92 0.73 0.51 0.91 0.86 0.91
(a) $Subject\ S1$
(b) $Subject\ S2$
(c) $Subject\ S3$
(d) $Subject\ S4$
(e) $Subject\ S5$
(f) $Subject\ S6$
Figure 3: Selected regression models across subjects based on the training set
VI Results
In this section, we report results pertaining to our research questions and discuss their practical implications.
VI-A RQ1: Can an accurate regression model be built to explain FDR as a function of MS?
To answer this question, we first investigate the extent of the correlation between MS and FDR. We perform our experiment on a large number of randomly and uniformly sampled input subsets from the training set, also considering different MS calculation variants. The correlation analysis results using Spearman correlation coefficients for experiment variants E1 and E2 are shown in Table III. We only report results for random subsets as they yielded slightly higher correlations and reporting results from uniform subsets does not provide new insights. The best correlation achieved for each subject across experiment variants, given in each column of Table III, is highlighted in bold font. We did not report the results for experiment variant E3 since all of the correlation values for both random and uniform subsets across all subjects are below 0.1, indicating an absence of correlation. This can be explained by the fact that MS is determined by computing the average $KillingScore$ (Equation (3)) of all inputs within a subset. Consequently, for subsets of different sizes, MS calculated based on $KSBasedMS$ (Equation (4)) falls within a similar range (0 to 0.1). As discussed in Section II-B, this confirms that calculating the MS of a subset as the average capability of its inputs to kill mutants cannot be used as an indicator of the capability of the entire subset to kill mutants.
E1 shows the strongest correlation across all subjects between MS and FDR (above 0.98 across all subjects). In this experiment, we calculate MS as proposed originally by DeepMutation (Equation (2)). In contrast, for experiment E2, $R^{2}$ ranges from 0.51 to 0.92, which is still a strong correlation. The method for identifying killed mutants and calculating MS therefore has a substantial impact on the results of mutation analysis. Given the strong correlations achieved in E1 and E2 across all subjects, we conclude that mutants generated by DeepMutation’s post-training MOs are highly associated with faults. Given the above results, we focus solely on $DeepMutationMS$ and random subsets from now on. Before proceeding to the next steps, we calculated the correlation between FDR and MS individually for each MO as well as for different subsets of operators. However, none of them achieved a significantly higher correlation than the correlation obtained with all MOs. Consequently, we proceeded to the next step considering, for each subject, all of the applicable MOs and performed regression analysis to model the relationship between MS and FDR on the training set. For each subject, we examined multiple regression models starting with a simple linear regression model and then using more complex non-linear models (quadratic, exponential, and regression trees) when needed. We then performed a 5-fold cross-validation procedure to calculate $R^{2}$, MMRE, and RMSE and selected the most accurate model for each subject. Due to space limitation, we only report the results for the selected regression model for each subject in Table IV. These results show a minimum $R^{2}$ value of 0.98 across all subjects, indicating that MS explains most of the variance in FDR. The average RMSE across all $K$-folds for all subjects ranges from 0.03 to 0.05, highlighting the high prediction accuracy. Additionally, the average MMRE varies between 0.1 and 0.18 indicating a low relative error. This suggests that MS can be used to very accurately predict FDR in the training set.
TABLE IV: Accuracy and shape of the selected regression models when predicting FDR with MS, based on the training set and using a 5-fold cross-validation procedure
Accuracy Subjects
metric S1 S2 S3 S4 S5 S6
$R^{2}$ 0.99 0.99 0.98 0.98 0.99 0.98
MMRE 0.18 0.13 0.15 0.14 0.1 0.15
RMSE 0.03 0.04 0.04 0.05 0.03 0.04
Shape* RT RT RT E RT RT
*Selected regression model shapes include Exponential (E) and Regression Tree (RT)
The final row of Table IV reports the shape of the selected regression model on each subject. Regression trees are the best for all subjects but S4. This suggests that standard regression functions are often not a good fit and regression trees offer more flexibility by optimally partitioning the MS range and computing averages within partitions. To prevent overfitting of the regression trees, which would yield poor results on test sets, we limited their maximum depth to 5.
To visualize our results, we present for each subject in Figure 3, the scatterplots and fitted regression curves between FDR and MS along with 95% prediction intervals (PI), based on all input subsets sampled from the training set. We compute PIs for regression tree models based on a bootstrapping method for non-parametric regression models [33]. PIs depict, for each specific MS value, the likely range of potential FDR values for a 95% confidence level. The narrow PIs across all subjects suggest that the regression models can be confidently used to predict the FDR of a test set with limited uncertainty. We can also observe that the relationship is clearly not linear in most cases. For all subjects, as subset size increases, MS initially shows rapid growth compared to FDR, indicating that the first mutants are relatively easy to kill when compared to faults. But then the rate of increase in MS starts to quickly slow down over time, suggesting that mutants are rapidly becoming more difficult to kill. This confirms that many mutants are not comparable to faults but that nevertheless does not prevent us from accurately modeling their relationship. The highly non-linear relationship between MS and FDR on most subjects indicates that MS cannot be directly used to assess test sets and decide on their adequacy. As illustrated in Figure 3, identical MS values can result in very different levels of fault detection abilities across subjects. For instance, in subject S6, an MS of 0.8 is associated with an FDR of 0.2 while in subject S4, the same MS is associated with an FDR of 0.8.
It should be noted that the shape of the relationship between MS and FDR greatly varies across subjects, which is likely due to the characteristics of the DNN models and associated training sets. For example, differences in internal architectures across DNN models determine the distribution of mutants generated by post-training MOs and contribute to such variation in relationships. Though this is not an issue in practice, this highlights the need to determine the best regression model empirically for each DNN model based on its training set, as specified by TEASMA (Section III).
Answer to RQ1: Post-training MOs generate mutants that are highly associated with faults; thus, based on these mutants, MS has a very strong correlation with FDR, especially when computed with $DeepMutationMS$. Further, an accurate regression model with a minimum $R^{2}=0.98$ (based on cross-validation) can be fitted for each subject based on its training set to accurately capture the relationship between FDR and MS. The shape of the regression model greatly varies across subjects.
VI-B RQ2: Can a test set’s FDR be accurately predicted from its MS using the best regression model built on the training set?
To address this question, we analyze the relationship of the predicted FDR ($\widehat{FDR}$) with the actual one ($ActualFDR$) in the test set. To do so, for each subject, we selected a large number of random test subsets and measured the fit ($R^{2}$) and slope of a linear regression model between $\widehat{FDR}$ and $ActualFDR$. If the former can be trusted as an indicator of the latter, we should have a high $R^{2}$ and a slope close to 1.
TABLE V: Linear correlation between $\widehat{FDR}$ and $ActualFDR$ of all random test subsets across subjects
Accuracy Subjects
metrics S1 S2 S3 S4 S5 S6
$R^{2}$ 0.97 0.98 0.94 0.97 0.98 0.95
Slope 1.21 1.19 1.16 1.23 0.95 1.15
Table V reports the results, where we can see a minimum $R^{2}$ value of 0.94 and a slope that is either slightly below or above 1, across all subjects. Such results indicate that the regression models developed on the training set yield FDR predictions that are closely aligned with the actual FDR of the test sets. Therefore, FDR predictions provide a good basis on which to assess the adequacy of test sets and whether their results can be trusted to decide about deploying a DNN model.
In Figure 4, we depict the linear regression lines between $\widehat{FDR}$ and $ActualFDR$ where the intercept is set to zero, clearly showing a strong alignment between the two. The slopes of the lines fall within the range [0.95 1.23], thus indicating no strong bias in the FDR predictions . In most subjects, the prediction models tend to slightly underestimate the actual FDR, with the exception of subject S5, where an overestimation is observed.
Answer to RQ2: $\widehat{FDR}$ is strongly correlated and therefore closely aligned with $ActualFDR$, thus indicating that, for each subject, the selected regression model built on the training set accurately predicts the FDR of a test set based on its MS. Consequently, TEASMA provides a reliable basis for assessing test sets and whether to trust their results when validating a DNN model.
(a) $Subject\ S1$
(b) $Subject\ S2$
(c) $Subject\ S3$
(d) $Subject\ S4$
(e) $Subject\ S5$
(f) $Subject\ S6$
Figure 4: Linear regression line between $\widehat{FDR}$ and $ActualFDR$
VI-C Threats to Validity
We discuss in this section the different threats to the validity of our results and describe how we mitigated them.
Internal threats to validity One of the internal threats to validity pertains to our sampling approach, as our results are based on sampled input subsets. To address this, we selected diverse input subsets, employing both random and uniform sampling techniques and varying sample sizes. Another internal threat to validity stems from selected parameters used throughout the experiments. This encompasses parameters for clustering mispredicted inputs to estimate faults and for configuring MOs. To mitigate this concern, regarding clustering parameters, we explored different settings for each subject and opted for the best configuration based on the evaluation of the Silhouette index [34] and DBCV scores [35], as suggested by Aghababaeyan et al. [6]. For MO configuration, we used the default settings of the DeepMutation++ tool, which yielded accurate results in terms of FDR prediction.
Construct threats to validity A potential threat is related to the calculation of MS and FDR. To mitigate this concern and ensure the robustness of our results, we considered multiple ways to compute MS, as described in Section II-B. Regarding FDR, we leveraged a SoTA fault estimation approach [6] based on clustering mispredicted inputs. This approach has been thoroughly validated and the resulting clusters have been carefully analyzed by the authors. We reused their publicly available approach to achieve accurate fault identification.
External threats to validity To address potential concerns regarding the generalizability of TEASMA, we performed our experiment on multiple subjects including various combinations of five DNN models with different internal architectures and four input sets containing diverse types of images. Another threat might be related to the specific set of MOs that we considered in our experiment. For practical reasons, we focused on post-training MOs and investigated all the MOs implemented by the DeepMutation++ tool, also considering whether subsets of MOs might yield better results. Other MOs may yield different results in the future.
VII The Cost of Using Pre-training Mutants
In this section, we compare the cost associated with performing mutation analysis based on pre-training mutants generated by DeepCrime with post-training mutants generated by DeepMutation++. Though this is not the primary focus of this paper, it provides initial evidence supporting our rationale for employing post-training MOs in TEASMA and our concerns regarding the impractical cost of pre-training mutants, as previously outlined in Section II-C. To perform this experiment we used the same computing environment described in Section IV-C. Given the experiment’s time-intensive nature, which entails consuming enormous amounts of computing resources, conducting this experiment on all our subjects would take many weeks and is neither feasible nor necessary for our purpose. This problem is made more acute by unpredictable server outages, which are unavoidable in such large experiments. As a result, we opted to focus on one of our smaller DNNs, specifically subject S1 (LeNet5), since it is expected to provide a conservative result regarding the cost of pre-training mutants while maintaining the computations of our experiment within a reasonable level. Indeed, the LeNet5 network architecture consists of seven layers containing basic deep learning modules including convolutional, average pooling, and fully connected layers. Consequently, the retraining process for generating each mutant is expected to be less time-consuming compared to deeper and more versatile network architectures such as ResNet20, which includes 20 layers. Therefore, performing the same experiment on subjects with larger DNNs and input sets is anticipated to require even more computational resources.
As described in Section II-C, the cost of mutation analysis is mainly due to performing two processes: mutant generation and MS calculation. TEASMA includes these two processes in its first step as described in Section III. Therefore, we repeated TEASMA’s first step utilizing DeepCrime’s pre-training MOs in place of DeepMutation++’s post-training MOs. Consequently, we can compare the cost and achieved Spearman correlation between FDR and MS based on mutants generated by DeepCrime and DeepMutation++ on identical randomly sampled input subsets of subject S1.
As described in Section II-B, when using pre-training MOs, different methods can be used for determining killed mutant and calculating MS. To conduct a comprehensive analysis we consider all the MS calculation variants for pre-training mutants detailed in Section II-B. As a result, we executed three variations of our cost analysis experiment, each differing in the way they calculate MS: (E4) $DeepCrimeMS$, (E5) $isKilledBasedMS$, and (E6) $DeepMutationMS$. Recall that E4 and E5 determine killed mutants based on Equation (5) while E6 does so based on Equation (9).
The duration of conducting mutation analysis using DeepMutation++ and DeepCrime on subject S1 is reported in Table VI. We must highlight that we adhere to the default procedure of each tool for filtering mutants with undesirable characteristics (equivalent, trivial, or redundant). The time taken for this procedure is included in the overall time of calculating MS. DeepCrime created 195 mutated versions of the training set or training program, retraining each version $20$ times (the tool’s default value), resulting in a total of 3900 mutant instances. As a result, Generating mutants using DeepCrime takes 43.52 hours. In contrast, DeepMutation++ directly mutated the original model and generated 304 mutated models in less than 10 seconds. Such a significant difference primarily results from the retraining process, and to some extent, is expected from any pre-training mutation testing tool. The overall time required for MS calculation based on pre-training mutants depends on the MS calculation approach and is 79.98 hours for E4 or E5 and 10.03 hours for E6. Both E4 and E5 use the $isKilled$ formula (Equation (5)) to determine killed mutants and calculate MS. In contrast, E6 follows the same MS calculation approach to DeepMutation++ (Equation (2)). Notably, the calculation of MS for the same set of input subsets, based on post-training mutants generated by DeepMutation++ and its way of calculating MS (E1) only takes 48.19 minutes.
TABLE VI: The total time of performing mutation analysis using DeepMutation++ and DeepCrime on subject S1
Tool DeepMutation++ DeepCrime
MS Formula E1 E4 or E5 E6
Mutant generation 9.10 s 43.52 h 43.52 h
MS calculation 48.19 m 79.98 h 10.03 h
Total 48.34 m 123.50 h 53.55 h
The time is reported in second (s), minutes (m), or hours (h)
The total time of performing mutation analysis based on the pre-training MOs of DeepCrime takes up to 123 hours when using DeepCrime’s statistical MS calculation approach (E4). This is approximately 153 times longer than the duration based on post-training MOs with DeepMutation++ and its original MS calculation (E1). Given the considerable time investment and the substantial computational resources needed for analyzing even a relatively small subject, the practicality of using pre-training MOs is limited in numerous application scenarios. Indeed, this total time is computed for the LeNet5 network, which is comparatively small compared to typical contemporary networks in industry. Utilizing a larger DNN network with more layers would increase mutant generation time due to the extended retraining time for each mutant. Similarly, larger input sets would contribute to an increased training time for each mutant. Consequently, as DNNs and input sets grow in size, the difference in the time required for performing mutation analysis based on pre- and post-training MOs is expected to increase.
To investigate whether the enormous cost of pre-training mutants would somehow be justified by improved FDR predictions, we also computed the Spearman correlation between FDR and MS based on E4, E5, and E6. This resulted in correlation values of 0.76, 0.61, and 0.98, respectively. We note that E4, representing the statistical approach proposed by DeepCrime for calculating MS, yields a much weaker correlation than E1 (0.74 vs. 0.99), the original way of calculating MS based on post-training mutants in DeepMutation++, as visible in Table III. Further, E6, which relies on the same formula as E1 but with pre-training MOs, shows a similar correlation to E1 with 0.98. Based on the above, we can conclude that there is no advantage, in terms of FDR prediction, to using pre-training MOs in DeepCrime. However, the execution of E6 required approximately 66 times the computation time of E1. Consequently, post-training MOs in DeepMutation++ clearly require much less computational resources for conducting mutation analysis. While this experiment, due to its high computational requirements, is conducted on a single subject of modest size, the findings underscore the substantial difference in computational resource requirements for mutation analysis based on pre- and post-training MOs. This observation, in addition to the strong association observed between mutants generated by post-training MOs and faults, further justifies the choice made in TEASMA to rely on post-training MOs.
VIII Conclusion
In this paper, we propose TEASMA, a practical approach for assessing the adequacy of test sets for DNNs. TEASMA is based on mutation analysis during the training phase. Such assessments are important in practice, regardless of the test set selection strategy, in order to determine whether test results can be trusted to validate DNN models.
Our results show that, although there is a strong relationship between Mutation Score (MS) and Fault Detection Rate (FDR), the shape of the relationship is typically not linear and greatly varies across DNN models. Consequently, MS is not an accurate surrogate measure for FDR and cannot be used directly as a test adequacy criterion. TEASMA’s strategy is therefore to provide a means to accurately predict the FDR of a test set from its MS so as to confidently decide whether it is adequate to provide reliable test results. If deemed adequate, the test set is then labeled and executed for validating the DNN model. Such a procedure also prevents the potential waste of effort associated with labeling test inputs that may turn out to be inadequate.
TEASMA relies on post-training mutation operators as they require much less computational resources compared to pre-training operators. Post-training operators do not require retraining and are therefore practically applicable for test adequacy assessment in all situations. Though not representative of real faults, our results show that such operators can be used for such assessments since we can accurately predict FDR based on MS.
TEASMA relies on regression analysis to build an accurate FDR prediction model based on MS using the training set. Based on training sets, prediction models yielded a very high $R^{2}$ ($R^{2}\geq 0.98$) across subjects and thus enabled TEASMA to confidently predict the FDR of an input set from its MS. Our results further indicate that such prediction models can accurately predict a test set’s FDR— with $R^{2}\geq 0.94$ and a slope close to 1—which can then reliably be used to decide about the adequacy of a test set.
According to our findings, TEASMA is thus a reliable instrument to ensure the proper validation of a DNN before deployment, especially in more critical contexts where trustworthiness is crucial.
In our future work, we intend to rely on TEASMA to automate the generation of adequate test sets for DNN models.
Acknowledgements
This work was supported by a research grant from Huawei Technologies Canada Co., Ltd, as well as by the Mitacs Accelerate Program. Additional support was provided by the NSERC Discovery and Canada Research Chair programs. The experiments conducted in this work were enabled in part by support provided by the Digital Research Alliance of Canada  [^1] https://alliancecan.ca.